OpenStudy (anonymous):

*HELP!** Factor this polynomial completely over the set of complex roots. f(x)=x^3-x^2-4x+4

OpenStudy (anonymous):

see this \( x^3-x^2-4x+4=x^2(x-1)-4(x-1) \)

OpenStudy (anonymous):

How many zeros?

OpenStudy (anonymous):

2 zeros

OpenStudy (anonymous):

\( x^3-x^2-4x+4=x^2(x-1)-4(x-1)=(x-1)(x^2-4)=(x-1)(x-2)(x+2) \)

OpenStudy (anonymous):

opps 3 zeros

OpenStudy (anonymous):

if @experimentX taught me right yesterday.

OpenStudy (experimentx):

oh ... that was just clearing up confusion ... lol

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

yeah